# Episodic Memory Verbalization using Hierarchical Representations of Life-Long Robot Experience

### v1

@@ IV-B Egocentric Human Videos
Next to verbalizing robot experience, EMV can be applied to human egocentric recordings, e. g., in the context of smart wearables. Here, the system does not summarize and answer questions about its own actions, but the actions of its user.
@@
Results:   The results of applying our method and manually categorizing the results (following Table I) can be seen on the left of Table II. Due to the very challenging nature of our QA samples and the limitations of the used vision components, the overall performance is low. Low performance of the Gemini 1-pass baseline can partially be explained by it seeing only a flat version of the L2 events without access to the images. However, it also fails on most of the samples that could be answered from the text history. This may be explained by the noisy text history inferred from vision, which H-Emv can handle better because of hierarchical summarization and selective expansion of nodes, whereas the 1-pass baseline observes all the noise at once. We did not directly apply the 1-pass baseline on the images (which would be possible only with aggressive sub-sampling) for cost reasons. H-Emv L3 performs similar to H-Emv, with double the costs.
@@ IV-C Real-World Robot Recordings
Finally, we apply our method on the real-world humanoid robot Armar-7. To obtain an EM, we record multiple robot sessions of typical household tasks, spanning a total duration of 3.3 hours of robot actions over the scope of two months. We record all entries made to the memory system introduced in [22], in particular: vision (RGB and depth images), robot state (proprioception), skill events (executed actions and goals), speech (speech-to-text output and text-to-speech input), symbolic scene (objects and their relations). From such recordings, we build up a history tree by populating L0 with images, speech, and proprioception, L1 scene graphs with the symbolic scene information, L2 and L3 with robot action events (where L2 contains low-level actions and L3 contains actions that themselves invoke other actions). Note that L3 nodes can be nested in this case (goals and their subgoals). Higher levels (L4+) are constructed dynamically by an LLM as described in Sec. III-A, with two manually created few-shot samples.
Subsequently, we annotate the recordings with 30 QA-pairs, apply our method, and again manually categorize the results. Results can be seen on the right part of Table II. In general, our task is very challenging, and the 1-pass Gemini baseline which has direct access to the complete stream of episodic data (without images) scores only 40%/10% of correct/partially correct samples. Compared to the Ego4D experiment, the quality of the text history is better, as most content (esp. current action, goal) is not inferred from vision. Our interactive hierarchical system achieves slightly better performance, with 1/17 of the token costs. The numbers also highlight that the hierarchical aspect is crucial, as H-Emv with only L3 has notably lower performance with more than 5 times the token cost. See the supplementary video for a demonstration of our system in action, enabling Armar-7 to answer questions about its past interactively.
@@
B R $S_{c}$ $S_{p}$ T B R $S_{c}$ $S_{p}$ T
H-Emv 1.1 11.6 28 25 21.9 7.0 18.6 43 27 12.8
H-Emv (L3) 0.9 13.7 25 18 57.6 2.5 20.4 30 17 68.4
Gemini 1-pass flat 0.8 4.7 5 8 438 12.4 28.8 40 10 227
TABLE II: Experiments on Human & Robotic Real-World Data (0-shot)
@@ V Conclusion & Discussion
We present H-Emv, a system for verbalization of life-long robot experience. The multimodal, hierarchical representation of EM is interactively accessed by an LLM to answer user questions, keeping token costs low even for extremely long histories. Despite the promising results and versatility of our system, it has some limitations: First, as a modulated approach, it is limited by the performance of each component and can suffer from error propagation. While the interactive tree search improves interpretability, there are no performance guarantees. Moreover, our system could integrate more modalities and tools. For instance, joint angle proprioception data could be rendered in simulation and then verbalized by a VLM. Adding personalization, both to EM and verbalization, is desirable for improved human-robot interactions. We hope our code and data will foster research on EMV, and will continue addressing these challenges in future work.

### v2

@@ IV-B Egocentric Human Videos
In addition to verbalizing robot experience, EMV can be applied to human egocentric recordings, e. g., in the context of smart wearables. Here, the system does not summarize and answer questions about its own actions, but the actions of its user.
@@
Results:   The results of applying our method and manually categorizing the results (following Fig. 4) can be seen on the left of Table II. Due to the very challenging nature of our QA samples and the limitations of the used vision components, the overall performance is low. Low performance of the 1-pass baseline can partially be explained by it seeing only a flat version of the L2 events without access to the images. However, it also fails on most of the samples that could be answered from the text history. This may be explained by the noisy text history inferred from vision, which H-Emv can handle better because of hierarchical summarization and selective expansion of nodes, whereas the 1-pass baseline observes all the noise at once. We did not directly apply the 1-pass baseline on the images (which would be possible only with aggressive sub-sampling) for cost reasons. H-Emv L3 performs slightly worse than H-Emv, with double the costs.
Figure 5: Example traces from our Armar-7 evaluation. Left: success case. Middle: partially correct, LLM misinterpreting the question. Right (shortened): wrong answer, expanded the wrong node (not the last time grasping the milk). More examples at hierarchical-emv.github.io
@@ IV-C Real-World Robot Recordings
Finally, we apply our method on our real-world humanoid robot Armar-7. To obtain an EM, we record multiple robot sessions of typical household tasks, spanning a total duration of 3.3 hours of robot actions over the scope of two months. We record all entries made to the memory system introduced in [24], in particular: vision (RGB and depth images), robot state (proprioception), skill events (executed actions and goals), speech (speech-to-text output and text-to-speech input), symbolic scene (objects and their relations). From such recordings, we build up a history tree by populating L0 with images, speech, and proprioception, L1 scene graphs with the symbolic scene information, L2 and L3 with robot action events (where L2 contains low-level actions and L3 contains actions that themselves invoke other actions). Note that L3 nodes can be nested in this case (goals and their subgoals). Higher levels (L4+) are constructed dynamically by an LLM as described in Sec. III-A, with two manually created few-shot samples.
TABLE II: Experiments on Human & Robotic Real-World Data (0-shot)
@@
Method Hierarchy LLM $S_{c}$ $S_{p}$ T $S_{c}$ $S_{p}$ T
1-pass flat Gemini 5 13 438 43 53 227
H-Emv L3 Gemini 18 38 33.1 40 70 29.2
H-Emv full Gemini 30 40 14.4 33 73 10.5
H-Emv L3 GPT-4o 25 43 57.6 30 47 68.4
H-Emv full GPT-4o 28 53 21.9 43 70 12.8
Subsequently, we annotate the recordings with 30 QA-pairs, apply our method, and again manually categorize the results. Results can be seen on the right part of Table II, with examples in Fig. 5. In general, our task is very challenging, and the 1-pass Gemini baseline which has direct access to the complete stream of episodic data (without images) scores only 43%/53% of correct/partially correct samples. Compared to the Ego4D experiment, the quality of the text history is better, as most content (esp. current action, goal) is not inferred from vision. Our interactive hierarchical system achieves better $S_{p}$ and slightly worse $S_{c}$, with 1/21 of the token costs. While the experiments with Gemini and limited tree depth (L3) perform slightly better, the best results are achieved by the full system with GPT-4o. Here, L3 has lower performance with more than five times the token cost. See the supplementary video for a demonstration of our system in action, enabling Armar-7 to answer questions about its past interactively. We also applied H-Emv (with no further dataset-specific prompt tuning) on NaVQA [28], but obtained an accuracy of only 32% on the descriptive questions, which we attribute to their dataset being focused on very specific details, whereas our method strives at tasks that require broader context and summarization capabilities.
IV-D Failure analysis
To better understand the potential for future improvement of our system, we analyze its failure modes. We broadly categorize failures by three sources: EM construction, EM retrieval, and LLM QA reasoning. For the simulated TEACh evaluation, full multimodal setting, both the generated QA and the history tree are derived from the simulation state, therefore the requested information is guaranteed to be contained in the tree, eliminating the EM construction failure reason. During QA generation, we also annotate the target episode this question asks for, and then evaluate whether our model correctly expands the history tree node representing this episode. Specifically, we calculate retrieval precision $p$ and recall $r$ as follows: Given the final state of the tree when the LLM decides to answer the question, we extract the set of visible (expanded) nodes $V$. From these, we pick the lowest-level visible nodes: $L=\{v\in V\ |\ \text{children}(v)\cap V=\varnothing\}$. Every node $v$ has an associated time span $t_{v}=(s_{v},e_{v})$ with start and end datetime $s_{v},e_{v}$. Given a reference time span $t_{g}=(s_{g},e_{g})$ that the question refers to (one question may refer to multiple time spans, e. g., “what did you do before…”), we calculate the length of the intersection $i_{g,v}=|t_{g}\cap t_{v}|$. This intersection is normalized relative to the duration of the node and the baseline intersection $b$ of the root node $v_{r}$ with the reference, defined as $b=\frac{|t_{g}|}{|t_{v_{r}}|}$. The normalized intersection is given by $n_{g,v}=\max(0,\frac{i_{g,v}}{|t_{v}|}-b)\cdot\frac{1}{1-b}$. Recall is defined as the average of the maximum normalized intersection rate per reference time span: $r=\text{avg}_{t_{g}\in G}(\ \max_{v\in V}n_{g,v}\ )$. Precision is calculated by taking the maximum normalized intersection rate for each leaf $l\in L$ over all reference spans, and then averaging across all leaves: $p=\text{avg}_{l\in L}(\ \max_{t_{g}\in G}n_{g,l}\ )$. Eventually, $F_{1}=\frac{p\cdot r}{p+r}$.
Fig. 6 shows the retrieval $F_{1}$ of H-Emv for different history lengths. Evidently, the retrieval process is a major source of problems for the EMV task: When taking into account only samples that were answered (partially) correct, the retrieval performance is notably better than for the samples that were answered wrong, underlining that correct retrieval is essential for overall QA performance. The non-zero retrieval performance for wrong samples stems from partial retrieval (not expanded deep enough) as well as other failures due to QA reasoning (despite correct retrieval).
5515152525505010010000.20.20.40.40.60.60.80.811|h||h|Retrieval $F_{1}$overallonly correctonly part. corr.only wrong
Figure 6: Retrieval $F_{1}$ score by history length $|h|$ (TEACh H-Emv 1-shot)
For our real-world (Ego4D and Armar-7) experiments, we manually inspect the execution traces of H-Emv (Gemini) for each (partially) wrong sample and attribute it with one or multiple of the above failure reasons. As for the simulated evaluation, the majority of failures is related to retrieval (18 of 28 failure samples for Ego4D, 19/20 for Armar-7). Especially for the real-robot experiments, the tree is very deeply nested, and retrieval therefore challenging. On the other hand, tree building issues are more prevalent for Ego4D (13/28), as the system is relying purely on visual information there. However, issues can also occur on Armar-7 (7/20), for instance when the LLM-generated summaries are misleading or too broad (“performed various actions”). Further, both approaches sometimes struggle with reasoning or correctly using the tree exploration API (e. g., getting stuck in repetitive behavior), which is most likely related to the difficulty of the questions and the general LLM capabilities.
@@ V Conclusion & Discussion
We present H-Emv, a system for verbalization of life-long robot experience. The multimodal, hierarchical representation of EM is interactively accessed by an LLM to answer user questions, keeping token costs low even for extremely long histories. Despite the promising results and versatility of our system, it has some limitations: First, as a modulated approach, it is limited by the performance of each component and can suffer from error propagation. While the interactive tree search improves interpretability, there are no performance guarantees. Moreover, our system could integrate more modalities and tools, e. g., joint angle proprioception data could be rendered in simulation and then verbalized by a VLM. Episodic memories should refer back to previous events to enable contextual or comparative descriptions [66]. Adding personalization, both to EM and verbalization, is desirable for improved human-robot interactions. We hope our code and data will foster research on EMV, and will continue addressing these challenges in future work.